# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 1. $(\mathbf{noc_{1}^{\prime}})$
Simply observe that the proof of Theorem 29 does not make use of the fact that quantum polymorphisms of $\mathbb{C}_{m}$ are oracular. Hence, the same argument establishes that $\operatorname{qPol^{no}}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$ for each odd $m\geq 3$. Then, the result follows from Theorem 38 applied to $\mathbb{A}=\mathbb{B}=\mathbb{C}_{m}$ and Corollary 34. ∎
8. Quantum polymorphisms of Boolean structures
In this section, we provide some results regarding the quantum polymorphisms of Boolean structures. If a quantum homomorphism $Q$ has range $\{0,1\}$, then $Q_{x,0}$ is determined by $Q_{x,1}$ since $Q_{x,0}=\operatorname{id}-Q_{x,1}$. Moreover, the set $\{0,1\}^{n}$ is naturally in bijection with the powerset of $\{1,\dots,n\}$. Given $S\subseteq[n]$, we let $s\in\{0,1\}^{n}$ be the tuple such that $s_{i}=1$ iff $i\in S$. In the following, if $\mathbb{A}$ is a Boolean structure, we write the projectors of a quantum homomorphism $\mathbb{A}^{n}\to_{q}\mathbb{A}$ by $Q_{S}$ where $S\subseteq[n]$.
8.1. The case of structures without a majority polymorphism
In this subsection we show that any Boolean relational structure which is not invariant under the majority polymorphism has a commutativity gadget. This specific statement has previously been proven in [culf_re-completeness_2025, Section 7], our contribution is to simplify the proof by using the characterization in Theorem 44 to assert the non-existence of commutativity gadgets through the study of quantum polymorphisms. This allows us to make use of standard CSP methods in the proof.
We focus our study on the following class of Boolean relational structures:
@@ 8.1. The case of structures without a majority polymorphism
Let $t\in\{0,1\}^{k}$ be an arbitrary Boolean tuple. Then the $t$-translate of 11-in-$k$ is the relational structure $\mathbb{O}_{t}=(\{0,1\};\{x\oplus_{2}t\mid x\in R_{1/k}\})$, where
  1. (1)
@@
  2. (2)
    $\oplus_{2}$ is componentwise addition modulo 2.
@@
One can easily see that the ternary majority function on $\{0,1\}$ does not preserve $R_{t}$; the majority of any 3 distinct tuples in the relation yields $t\not\in R_{t}$, witnessing the non-preservation.
@@
For every non-oracular polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$, the following properties hold:
  1. (1)
@@
In particular, it follows that in every non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\to_{q}\mathbb{O}_{(1,0,0)}$ and every $S,T\subseteq[n]$, we have $Q_{S},Q_{T}]=0$, which implies that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})$ is the quantum closure of $\operatorname{Pol}(\mathbb{O}_{(1,0,0)})$.
@@
Since $(0,0,\dots,0)=(1,0,\dots,0)\oplus_{2}(1,0,\dots,0)\in R_{(1,0,\dots,0)}$ and $(1,1,\dots,1)\notin R_{(1,0,\dots,0)}$, we can pp-define 0 in $\mathbb{O}_{(1,0,\dots,0)}$ by
@@
Now assume we have some non-oracular quantum polymorphism $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\to_{q}\mathbb{O}_{(1,0,\dots,0)}$ and consider:
@@
$=(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})(Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})$
@@
Assume $k\geq 3$, and neither $t$ nor $\overline{t}$ ($(1,\dots,1)\oplus_{2}t$) are in $R_{1/k}$. Then $\operatorname{qPol^{no}}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$.
@@
If $l\notin\{1,k-1\}$ (i.e., neither $t$ nor $\overline{t}$ are in $R_{1/k}$) then $(0,\dots,0)=t\oplus_{2}t$ and $(1,\dots,1)=t\oplus_{2}\overline{t}$ are not in $R_{t}$. We first pp-define $\neq$ in $\mathbb{O}_{t}$ via
@@
where we fill the first $l+1$ entries with $x$’s. This relation is preserved by every non-oracular $Q\colon\mathbb{O}_{t}^{n}\to\mathbb{O}_{t}^{n}$ since for every $\mathbf{a}\neq\mathbf{b}\in\{0,1\}^{n}$ and $x\in\{0,1\}$ we have
@@
Now let $Q\colon\mathbb{O}_{t}^{n}\to_{q}\mathbb{O}_{t}$ be a non-oracular polymorphism of $\mathbb{O}_{t}$. Then we have
@@
Next, we argue that the only real cases we need to consider in the full theorem is the $\mathbb{O}_{t}$’s.
Lemma 54.
Every relational structure $\mathbb{A}$ with a relation $R$ that has any full binary projection has the property that $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
Proof.
Consider a relational structure with a relation $R$ that has a full binary projection on the components $\{i,j\}$. If this is the case, then in any power, any two elements appear together in some tuple of the direct power of RR.∎
@@
2. implies 1. Note that translating a relation by a tuple does not change the size of the binary projections and does not change preservation by majority, since majority is a self-dual operation. It follows that it suffices to prove the implication in the case of $R_{1/k}$, for which all the properties in 1. are clear.
For 1. implies 2. suppose $R\subseteq\{0,1\}^{k}$ is not invariant under majority but majority does preserve every proper projection and no binary projection is full. Suppose $t\in\{0,1\}^{k}$ is a witness to the fact that $R$ is not invariant under majority, we first show that $\{x\oplus_{2}t\mid x\in R_{1/k}\}\subseteq R$.
For $i\in[k]$, let $x\in R_{1/k}$ be the tuple with a 1 in the $i$th position. We know that $\pi_{[k]\setminus\{i\}}(R)$ is invariant under majority so there is an $r_{i}\in R$ with $\pi_{[k]\setminus\{i\}}(r_{i})=\pi_{[k]\setminus\{i\}}(t)$. Now, since $t\notin R$; we can conclude that $\pi_{\{i\}}(r_{i})\neq\pi_{\{i\}}(t)$ and therefore $x\oplus_{2}t=r_{i}$, so $x\oplus_{2}t\in R$.
Finally, we use the non-fullness of binary projections to argue that any $r\in R$ must match $\overline{t}=(1,\dots,1)\oplus_{2}t$ in exactly one argument, placing it in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$.
Assume that $r$ matches $\overline{t}$ in two arguments $1\leq i<j\leq k$. We know that $\pi_{i,j}(R)$ contains $\pi_{i,j}(t)$, $(\pi_{i}(t),\pi_{j}(\overline{t}))$ and $(\pi_{i}(\overline{t}),\pi_{j}(t))$; then $\pi_{i,j}(R)$ is full, a contradiction. Additionally, $r$ cannot match $\overline{t}$ on zero arguments since then it would equal $t$, which is not in $R$. Thus every $r$ is in $\{x\oplus_{2}t\mid x\in R_{1/k}\}$, completing the proof. ∎
We can now easily prove the general result.
@@
If $\mathbb{A}$ is a Boolean relational structure such that $\operatorname{Pol}(\mathbb{A})$ does not contain majority, then $\operatorname{qPol^{no}}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
@@
Consider a relational structure $\mathbb{A}$; without loss of generality, every projection of a relation of $\mathbb{A}$ is also a relation of $\mathbb{A}$, since this does not change the hypothesis or the conclusion of the statement. Let $R$ be a relation of $\mathbb{A}$ that is not preserved by majority and of minimal arity. By Lemma 54, if a binary projection of $R$ is full then we are done, so assume that this is not the case. Then Theorem 55 implies that $R=R_{t}$ for some $k$-tuple $t$.
Finally, we know that $k\geq 3$ since every binary relation of arity 2 is preserved under majority and so, most of the result follows from Corollary 52 and Lemma 53. All that remains is for us to argue that Lemma 52 also covers the case where $\overline{t}\in R_{1/k}$.
This is since, $\mathbb{O}_{t}$ has ‘essentially’ the same polymorphisms as $\mathbb{O}_{\overline{t}}$ for any $t$. The formal fact we use is that $\{Q_{S}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{t}$ if and only if $\{\operatorname{id}-Q_{\overline{S}}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{\overline{t}}$. In particular $\mathbb{O}_{t}$ has contextual polymorphisms if and only if $\mathbb{O}_{\overline{t}}$ does. Thus we can assume that $\overline{t}\notin R_{1/k}$, completing the proof. ∎
8.2. The minimal contextual clone
In this section, we provide a clone with contextual polymorphisms that is contained within the quantum polymorphism clone of any Boolean relational structure with contextual polymorphisms.
In the previous section, we have seen that every boolean relational structure with contextual polymorphisms has non-full binary projections and contains majority in its (classical) polymorphism clone.
Let $\mathbb{B}$ be the structure $\mathbb{B}:=(\{0,1\};S_{00},S_{11},S_{10})$ where $S_{ab}=\{0,1\}^{2}\setminus\{(a,b)\}$. In the classical setting, $\operatorname{Pol}(\mathbb{B})$ is the smallest clone that contains the majority operation. We first show that this property extends to oracular quantum polymorphisms.
Lemma 57.
Let $\mathbb{A}$ be a Boolean structure that admits majority as a polymorphism, and such that no binary projection of a relation of $\mathbb{A}$ is full. Then $\operatorname{qPol}(\mathbb{B})\subseteq\operatorname{qPol}(\mathbb{A})$.

### v2

@@ 1. $(\mathbf{noc_{1}^{\prime}})$
Let $g_{1},\dots,g_{r}$ be the distinguished vertices of $\mathbb{G}$. We only need to prove $($$\mathbf{QH_{1}}$)). Let $(a_{1},\dots,a_{r})\in R_{A}$ and $(b_{1},\dots,b_{r})\not\in R_{B}$. By ($(\mathbf{noq_{1}})$), there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a_{i}}=\operatorname{id}$ for all $i\in[r]$. Let $S=Q\bullet Q^{\prime}$, which is a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{B}$. By ($(\mathbf{noq_{2}})$), we have that the quantum function $Q^{\prime\prime}$ defined by $Q^{\prime\prime}_{i,b}:=S_{g_{i},b}$ for all $i\in[r]$ and all $b\in B$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}(B;R_{B})$. Thus, $\prod_{i=1}^{r}Q^{\prime\prime}_{i,b_{i}}=0$ and therefore $\prod_{i=1}^{r}S_{g_{i},b_{i}}=0$. Note that $S_{g_{i},b_{i}}$ is by definition $\sum_{a\in A}Q_{a,b_{i}}\otimes Q^{\prime}_{g_{i},a}=Q_{a_{i},b_{i}}\otimes\operatorname{id}$. Thus, $(\prod_{i=1}^{r}Q_{a_{i},b_{i}})\otimes\operatorname{id}=0$, from which we obtain that $\prod_{i=1}^{r}Q_{a_{i},b_{i}}=0$. ∎
Remark 102.
Let $Q\colon\mathbb{X}\operatorname{\to_{\operatorname{qno}}}\mathbb{Y}$ be a non-oracular quantum homomorphism between two digraphs $\mathbb{X}$ and $\mathbb{Y}$. Suppose that $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ are such that there exists a directed $\ell$-walk from $x$ to $x^{\prime}$ in $\mathbb{X}$ but there is no directed $\ell$-walk from $y$ to $y^{\prime}$ in $\mathbb{Y}$. Consider the structure $\mathbb{G}$ consisting of a path of length $\ell$, and call $u,v$ the two endpoints of this path. Since this structure is a tree, Proposition 100 implies that $(\mathbb{G},u,v)$ is a q-no-definition for certain binary relations $R_{X}\subseteq X^{2},R_{Y}\subseteq Y^{2}$ in $\mathbb{X}$ and $\mathbb{Y}$, respectively. It follows from Proposition 101 that $Q$ is a non-oracular quantum homomorphism $(X;R_{X})\operatorname{\to_{\operatorname{qno}}}(Y;R_{Y})$. By assumption, $R_{X}$ contains $(x,x^{\prime})$ and $R_{Y}$ does not contain $(y,y^{\prime})$, and therefore $Q_{x,y}Q_{x^{\prime},y^{\prime}}=0$ by $($$\mathbf{QH_{1}}$)). Hence, Proposition 100 provides an alternative proof of [59, Lemma 4.12].
We now show that our undecidability result for odd cycles, Theorem 9, also holds in the non-oracular setting, by virtue of a result from [34].
Proposition 103.
For every odd $m\geq 3$, $\operatorname{CSP_{q}^{no}}(\mathbb{C}_{m})$ is undecidable.
Proof.
It was observed in [34, Proposition 5.27 & Lemma 4.9] that, if an undirected graph $\mathbb{A}$ having no 4-cycle admits a commutativity gadget, then it also admits a non-oracular commutativity gadget. (The result is phrased therein in terms of algebraic commutativity gadgets; however, as noted in Theorem 51 and Theorem 97, this is equivalent to our definitions.) Hence, Theorem 85 directly implies that $\mathbb{C}_{m}$ has a non-oracular commutativity gadget. Note now that $\mathbb{C}_{m}$ trivially pp-defines itself via a tree gadget. Hence, Proposition 100 implies that $\mathbb{C}_{m}$ q-no-defines itself. Using Proposition 98, we deduce that $\operatorname{CSP_{q}}(\mathbb{C}_{m})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{C}_{m})$ in logspace. Hence, it follows from Theorem 9 that $\operatorname{CSP_{q}^{no}}(\mathbb{C}_{m})$ is undecidable. ∎
We conclude this section by observing that the notion of q-no-definitions can be extended to a notion of q-no-constructions in a similar way as in the oracular case considered in Section 5.1.
@@
Let $\mathbb{A}$ be a $\sigma$-structure and let $\mathbb{B}$ be a $\rho$-struture. We say that $\mathbb{A}$ q-no-constructs $\mathbb{B}$ if there exists a $\rho$-structure $\mathbb{C}$ such that
  • $(i)$
    the domain of $\mathbb{C}$ is $A^{d}$ for some $d\in\mathbb{N}$;
  • $(ii)$
    each relation of $\mathbb{C}$ (of arity, say, $r$) is q-no-definable in $\mathbb{A}$ (identifying $C^{r}=(A^{d})^{r}$ with $A^{dr}$);
  • $(iii)$
    $\mathbb{B}\to_{\operatorname{q}}\mathbb{C}$ and $\mathbb{C}\to_{\operatorname{q}}\mathbb{B}$.
Note that, in part $(iii)$, we require $\mathbb{B}$ and $\mathbb{C}$ to be quantum oracular homomorphically equivalent. As we shall see, non-oracular homomorphic equivalence is not sufficient in order to obtain a q-no-construction version of Proposition 98.
Proposition 105.
Let $\mathbb{A}$ and $\mathbb{B}$ be structures (with potentially different domains and signatures), and suppose that
  • $\mathbb{A}$ has a non-oracular commutativity gadget, and
  • $\mathbb{A}$ q-no-constructs $\mathbb{B}$.
Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace.
Proof.
Let $\sigma$ and $\rho$ be the signatures of $\mathbb{A}$ and $\mathbb{B}$, respectively, and let $\mathbb{C}$ be the $\rho$-structure with domain $A^{d}$ witnessing that $\mathbb{A}$ q-no-constructs $\mathbb{B}$, as per Definition 104. We now construct a structure $\mathbb{D}$ in the same way as in the proof of Theorem 4 in Section 5.1.
Given a tuple $\mathbf{a}=((a_{1,1},\dots,a_{1,d}),\dots,(a_{r,1},\dots,a_{r,d}))\in(A^{d})^{r}$, let $\bar{\mathbf{a}}\in A^{dr}$ be the tuple defined by
$\bar{\mathbf{a}}=(a_{1,1},\dots,a_{1,d},\dots,a_{r,1},\dots,a_{r,d}).$
Consider the signature $\tau$ having the same symbols as $\rho$ such that, if a symbol has arity $r$ in $\rho$, it has arity $dr$ in $\tau$. We let $\mathbb{D}$ be the $\tau$-structure with domain $D=A$ and relations defined as follows: For any $R\in\rho$, $R^{\mathbb{D}}=\{\bar{\mathbf{c}}:\mathbf{c}\in R^{\mathbb{C}}\}$.
Using part $(ii)$ of Definition 104, we know that $\mathbb{A}$ q-no-defines $\mathbb{D}$. Since, by assumption, $\mathbb{A}$ has a non-oracular commutativity gadget, we deduce from Proposition 98 that $\operatorname{CSP_{q}}(\mathbb{D})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace. On the other hand, we know from the proof of Theorem 4 that $\operatorname{qPol}(\mathbb{D})\to\operatorname{qPol}(\mathbb{C})$ and thus, by Theorem 3, $\operatorname{CSP_{q}}(\mathbb{C})$ reduces to $\operatorname{CSP_{q}}(\mathbb{D})$ in logspace. In addition, by part $(iii)$ of Definition 104, $\operatorname{CSP_{q}}(\mathbb{B})$ and $\operatorname{CSP_{q}}(\mathbb{C})$ are trivially interreducible in logspace. Combining the three reductions yields the required logspace reduction from $\operatorname{CSP_{q}}(\mathbb{B})$ to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$. ∎
Remark 106.
We point out that the Galois connection for q-no-definitions of Theorem 13 does not appear to lift to q-no-constructions. In particular, Theorem 4 does not immediately extends to the non-oracular setting. The reason is that the order we choose while identifying the elements of $A^{dr}$ with elements of $C^{r}$ (see the proof of Proposition 105) is now important, as in non-oracular homomorphisms we are not allowed to permute projectors associated with the same constraint. This, in turns, prevents from building a minion homomorphism $\operatorname{qPol^{no}}(\mathbb{D})\to\operatorname{qPol^{no}}(\mathbb{C})$, which is needed in order to conclude that q-no-constructions yield minion homomorphisms between the minions of quantum non-oracular polymorphisms.
11 Quantum polymorphisms of Boolean structures
In this section, we consider the case of quantum polymorphisms of Boolean structures, and we prove the following result.
See 11
As a direct consequence of this result, we obtain a $\mathsf{P}$ vs. undecidable complexity dichotomy for quantum CSPs parameterised by Boolean structures. We let $\mathbb{XOR}$ be the Boolean structure having two ternary relation symbols $R_{+}$ and $R_{-}$, whose interpretations are
$R_{+}^{\mathbb{XOR}}$ $=\{(1,0,0),(0,1,0),(0,0,1),(1,1,1)\},$
$R_{-}^{\mathbb{XOR}}$ $=\{(0,1,1),(1,0,1),(1,1,0),(0,0,0)\}.$
Note that $\operatorname{CSP}(\mathbb{XOR})$ encodes the satisfiability problem for systems of linear equations mod $2$.
See 12
Proof.
If $\mathbb{A}$ does not pp-define $\mathbb{XOR}$, it was shown in [2] that $\operatorname{CSP_{q}}(\mathbb{A})=\operatorname{CSP}(\mathbb{A})$, and the latter problem is solvable in polynomial time via the bounded-width algorithm [8]. Suppose now that $\mathbb{A}$ $p$-defines $\mathbb{XOR}$. This implies that $\operatorname{Pol}(\mathbb{A})\subseteq\operatorname{Pol}(\mathbb{XOR})$. Since $\mathbb{XOR}$ is not preserved by majority, it follows that $\mathbb{A}$ has no majority polymorphism. Then, it follows from Theorem 11 that $\mathbb{A}$ admits a commutativity gadget. Applying Proposition 49, we deduce that $\mathbb{A}$ q-defines $\mathbb{XOR}$. This implies via Proposition 42 that $\operatorname{CSP_{q}}(\mathbb{XOR})$ reduces to $\operatorname{CSP_{q}}(\mathbb{A})$ in logspace. Since $\operatorname{CSP_{q}}(\mathbb{XOR})$ is undecidable (as proved in [67]), we conclude that $\operatorname{CSP_{q}}(\mathbb{A})$ is undecidable. ∎
Remark 107.
Corollary 12 should be compared with [63, Theorem 5.11 (a)]. The result established therein concerns Boolean constraint systems with contexts, where each context is a subset of variables of the CSP instance whose associated projectors are required to commute, see [63, Definition 3.1]. By modifying the contexts in the reductions used to prove undecidability, classical pp-definitions can be employed to obtain sound and complete reductions between such problems. Syntactically, this can be enforced by requiring that the language admits a full binary relation, see for example [2, §6.1]. To extend the result to arbitrary Boolean languages, the use of commutativity gadgets appears to be necessary; see also [33]. On the other hand, we point out that the dichotomy for Boolean constraint systems with contexts in [63] is incomparable to ours, in that it also addresses other variants of quantum satisfiability. In contrast, Corollary 12 applies only to finite-dimensional quantum strategies in the tensor-product model (see also the discussion in Section 2.8).
11.1 The case of structures without majority polymorphism
In this subsection, we study the quantum polymorphism minions of Boolean structures that are not preserved under majority. As a consequence, we give an alternative proof of the fact, originally proved in [33, Section 7], that any Boolean relational structure that is not invariant under the majority polymorphism has a commutativity gadget.
If a quantum homomorphism $Q$ has range $\{0,1\}$, then $Q_{x,0}$ is determined by $Q_{x,1}$ since $Q_{x,0}=\operatorname{id}-Q_{x,1}$. Moreover, the set $\{0,1\}^{n}$ is naturally in bijection with the powerset of $[n]$. Given $S\subseteq[n]$, we let $s\in\{0,1\}^{n}$ be the tuple such that $s_{i}=1$ if, and only if, $i\in S$. In the following, if $\mathbb{A}$ is a Boolean structure, we write the projectors of a quantum homomorphism $\mathbb{A}^{n}\to_{\operatorname{q}}\mathbb{A}$ by $Q_{S}$ where $S\subseteq[n]$.
We begin our analysis by focussing on the following class of Boolean structures.
Definition 108.
@@ 11.1 The case of structures without majority polymorphism
Let $t\in\{0,1\}^{k}$ be an arbitrary Boolean tuple. Then the $t$-translate of 11-in-$k$ is the relational structure $\mathbb{O}_{t}=(\{0,1\};\{x+_{2}t\mid x\in R_{1/k}\})$, where
  1. 1.
@@
  2. 2.
    $_{2}$ is componentwise addition modulo 2.
@@
One can easily see that the ternary majority function on $\{0,1\}$ does not preserve $R_{t}$; the majority of any three distinct tuples in the relation yields some $t$ that is not in $R_{t}$, witnessing the non-preservation.
@@
For every non-oracular polymorphism $Q\colon\mathbb{O}_{(1,0,0)}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{(1,0,0)}$, the following properties hold:
  1. 1.
@@
In particular, it follows that for every non-oracular quantum polymorphism
$Q\colon\mathbb{O}_{(1,0,0)}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{(1,0,0)}$
and every $S,T\subseteq[n]$, we have $Q_{S},Q_{T}]=0$, which implies that $\operatorname{qPol^{no}}(\mathbb{O}_{(1,0,0)})$ is the quantum closure of $\operatorname{Pol}(\mathbb{O}_{(1,0,0)})$.
@@
Since $(0,0,\dots,0)=(1,0,\dots,0)+_{2}(1,0,\dots,0)\in R_{(1,0,\dots,0)}$ and $(1,1,\dots,1)\notin R_{(1,0,\dots,0)}$, we can pp-define 0 in $\mathbb{O}_{(1,0,\dots,0)}$ by
@@
Now let $Q\colon\mathbb{O}_{(1,0,\dots,0)}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{(1,0,\dots,0)}$ be a non-oracular quantum polymorphism and consider:
@@
$=(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})(Q_{\mathbf{0},0}\dots Q_{\mathbf{0},0})$
$+(Q_{\mathbf{a},x}Q_{\mathbf{b},y}Q_{\mathbf{c},z}Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})(Q_{\mathbf{0},1}\dots Q_{\mathbf{0},1})$
@@
Assume $k\geq 3$, and neither $t$ nor $\overline{t}$ ($(1,\dots,1)+_{2}t$) are in $R_{1/k}$. Then $\operatorname{qPol^{no}}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$.
@@
If $l\notin\{1,k-1\}$ (i.e., neither $t$ nor $\overline{t}$ are in $R_{1/k}$) then $(0,\dots,0)=t+_{2}t$ and $(1,\dots,1)=t+_{2}\overline{t}$ are not in $R_{t}$. We first pp-define $\neq$ in $\mathbb{O}_{t}$ via
@@
where we fill the first $l+1$ entries with $x$’s. This relation is preserved by every non-oracular $Q\colon\mathbb{O}_{t}^{n}\to\mathbb{O}_{t}$ since for every $\mathbf{a}\neq\mathbf{b}\in\{0,1\}^{n}$ and $x\in\{0,1\}$ we have
@@
Now let $Q\colon\mathbb{O}_{t}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{O}_{t}$ be a non-oracular polymorphism of $\mathbb{O}_{t}$. Then we have
@@
Thus $\operatorname{qPol}(\mathbb{O}_{t})=\operatorname{\oplus Pol}(\mathbb{O}_{t})$ for every $k$-tuple $t$. We now argue that this result covers every Boolean structure that is of interest to us. Specifically, we establish the following result, using similar ideas as in [33, Proposition 7.19].
@@
  • $2.\Rightarrow 1.$
    Note that translating a relation by a tuple does not change the size of the binary projections and does not change preservation by majority, since majority is a self-dual operation. It follows that it suffices to prove the implication in the case of $R_{1/k}$, for which all the properties in $1.$ are clear.
  • $1.\Rightarrow 2.$
    Suppose $R\subseteq\{0,1\}^{k}$ is not invariant under majority but majority does preserve every proper projection and no binary projection is full. Suppose $t\in\{0,1\}^{k}$ is a witness to the fact that $R$ is not invariant under majority. We first show that $\{x+_{2}t\mid x\in R_{1/k}\}\subseteq R$.
    For $i\in[k]$, let $x\in R_{1/k}$ be the tuple with a 1 in the $i$th position. We know that $\pi_{[k]\setminus\{i\}}(R)$ is invariant under majority, so there is an $r_{i}\in R$ with $\pi_{[k]\setminus\{i\}}(r_{i})=\pi_{[k]\setminus\{i\}}(t)$. Now, since $t\notin R$, we can conclude that $\pi_{\{i\}}(r_{i})\neq\pi_{\{i\}}(t)$ and therefore $x+_{2}t=r_{i}$, so $x+_{2}t\in R$.
    Finally, we use the non-fullness of binary projections to argue that any $r\in R$ must match $\overline{t}:=(1,\dots,1)+_{2}t$ in exactly one argument, placing it in $\{x+_{2}t\mid x\in R_{1/k}\}$.
    Assume that $r$ matches $\overline{t}$ in two arguments $1\leq i<j\leq k$. We know that $\pi_{i,j}(R)$ contains $\pi_{i,j}(t)$, $(\pi_{i}(t),\pi_{j}(\overline{t}))$, and $(\pi_{i}(\overline{t}),\pi_{j}(t))$; then $\pi_{i,j}(R)$ is full, a contradiction. Additionally, $r$ cannot match $\overline{t}$ on zero arguments since then it would equal $t$, which is not in $R$. Thus every $r$ is in $\{x+_{2}t\mid x\in R_{1/k}\}$, completing the proof.∎
Since a structure is TVF if, and only if it doesn’t have any full binary projections, Proposition 61 allows us to apply Theorem 113 and prove the following:
@@
If $\mathbb{A}$ is a Boolean relational structure such that $\operatorname{Pol}(\mathbb{A})$ does not contain majority, then $\operatorname{qPol}(\mathbb{A})=\operatorname{\oplus Pol}(\mathbb{A})$.
@@
Consider a relational structure $\mathbb{A}$; without loss of generality, every projection of a relation of $\mathbb{A}$ is also a relation of $\mathbb{A}$, since this does not change the hypothesis or the conclusion of the statement. Let $R$ be a relation of $\mathbb{A}$ that is not preserved by majority and of minimal arity. By Proposition 61, if a binary projection of $R$ is full then we are done, so assume that this is not the case. Then Theorem 113 implies that $R=R_{t}$ for some $k$-tuple $t$.
Finally, we know that $k\geq 3$ since every binary relation of arity 2 is preserved under majority and so, most of the result follows from Corollary 111 and Lemma 112. All that remains is for us to argue that Corollary 111 also covers the case where $\overline{t}\in R_{1/k}$.
This is since, $\mathbb{O}_{t}$ has ‘essentially’ the same polymorphisms as $\mathbb{O}_{\overline{t}}$ for any $t$. The formal fact we use is that $\{Q_{S}\mid S\subseteq[n]\}$ is a non-oracular quantum polymorphism of $\mathbb{O}_{t}$ if, and only if, $\{\operatorname{id}-Q_{\overline{S}}\mid S\subseteq[n]\}$ is a (non-oracular) quantum polymorphism of $\mathbb{O}_{\overline{t}}$. In particular $\mathbb{O}_{t}$ has contextual polymorphisms if, and only if, $\mathbb{O}_{\overline{t}}$ does. Thus we can assume that $\overline{t}\notin R_{1/k}$, completing the proof. ∎
11.2 The minimal contextual minion
In this section, we provide a minion with contextual polymorphisms that is contained within the quantum polymorphisms of any Boolean relational structure that has contextual polymorphisms.
In the previous section, we have seen that every Boolean relational structure with contextual polymorphisms has non-full binary projections (is TVF) and contains majority in the minion of (classical) polymorphisms.
Let $\operatorname{\mathbb{M}}$ be the structure $\operatorname{\mathbb{M}}:=(\{0,1\};S_{00},S_{11},S_{10})$ where $S_{ab}=\{0,1\}^{2}\setminus\{(a,b)\}$. In the classical setting, $\operatorname{Pol}(\operatorname{\mathbb{M}})$ is the minimal clone containing the majority operation. We first show that this property extends to oracular quantum polymorphisms.
Theorem 115.
Let $\mathbb{A}$ be a Boolean TVF structure that admits majority as a polymorphism. Then $\operatorname{qPol}(\operatorname{\mathbb{M}})\subseteq\operatorname{qPol}(\mathbb{A})$.